Kate is 3 times as old as her son. If Kate is 30 years older than her son, what are their ages?
Answer by Cromlix(4381)

Hi there,
Kate is 3 times as old as her son.
K = 3S.....(1)
If Kate is 30 years older than her son
K = 30 + S.......(20
Replace K in Equation (2) with
3S from Equation (1)
3S = 30 + S
Collect like terms:
3S - S = 30
2S = 30
S = 15
Kate is currently 45 years old.
Her son is currently 15 years old.
Hope this helps :-)